### v1

# Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

Abstract
We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms. MAPF focuses on planning collision-free paths for a group of agents. While state-of-the-art MAPF algorithms can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear. Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms. Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings. SMART fills this gap with several advantages: (1) SMART uses a physics-engine-based simulator to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots. In addition, we use SMART to explore and demonstrate research questions about the execution of MAPF algorithms in real-world scenarios. The code is publicly available at https://jingtianyan.github.io/publication/2025-smart.
1 Introduction
Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c) Robot model used in the environment. (d) A 3D view of the environment.
The Multi-Agent Path Finding (MAPF) problem Stern et al. (2019) focuses on planning collision-free paths for large groups of robots within a known environment. State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds Ma et al. (2019); Li et al. (2021c); Zhang et al. (2022); Li et al. (2022). However, these algorithms make several simplifying assumptions. First, they often rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths. Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors. Third, they typically plan and evaluate paths on unrealistic discretized 2-D grid maps Stern et al. (2019). Therefore, while researchers have shown that MAPF algorithms have a wide range of applications Ma et al. (2017); Hönig et al. (2019); Ho et al. (2022); Li et al. (2023); Zhang et al. (2023b, c, 2024); Zang et al. (2025), the algorithms are mostly evaluated under simplified settings, leaving their performance in more realistic settings unclear.
MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings. In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF algorithms is impractical. Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers. Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans. Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort. While some software tools Schaefer et al. (2023); Heuer et al. (2024) have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps. In addition, industrial professionals possess realistic simulators for their applications of interest. However, they may not have the resources and expertise to implement state-of-the-art MAPF algorithms in their simulators. By having a realistic simulator that can easily interface with existing implementations of MAPF algorithms, industrial professionals can more easily understand and select a MAPF algorithm that fits their demands.
In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios. SMART comprises three main components: (1) a physics-engine-based simulator based on Argos 3 Pinciroli et al. (2012), (2) an execution monitoring (EM) server, and (3) robot-specific executors. The simulator generates realistic simulation environments based on user-defined configurations. Fig. 1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark Stern et al. (2019). The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors. It leverages the Action Dependency Graph (ADG) Hönig et al. (2019), a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models. Each robot is assigned an individual executor that receives and executes actions from the server. SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and spatial inaccuracies caused by controller limitations.
Our contributions in this paper are: (1) We present SMART, an open-source software tool to evaluate MAPF algorithms in realistic scenarios, which can scale to thousands of robots while ensuring replicability, (2) we evaluate SMART with a variety of open-source implementations of MAPF algorithms and robot models, showing its flexibility, and (3) we show some important findings regarding the application of MAPF algorithms in realistic scenarios. We also use them as examples of how SMART can be used to address such problems.
Figure 2: (a) MAPF problem. (b) Associated MAPF plan. (c) ADG from the MAPF plan. Green ovals are actions, with the first element as the reference time from the plan, the second as the action type, and the last two as start and goal locations. Horizontal edges are Type-1, and vertical edges are Type-2. All robots initially face north.
Figure 3: System overview of SMART.
2 Background
2.1 MAPF Problem and MAPF Plan
A MAPF instance Stern et al. (2019) consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ and a set of $I$ robots =$\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations. Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations. Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex. Two robots collide if they are at the same vertex or swap vertices at the same timestep. An example MAPF instance and plan on a 2-D grid graph is shown in Fig. 2 (a) and (b), respectively. The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
Prior research has proposed many MAPF algorithms. Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees Sharon et al. (2015); Barer et al. (2014); Li et al. (2021c), while others focus on scalability, solving MAPF problems involving thousands of robots Okumura et al. (2022); Okumura (2023). Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency Li et al. (2021a); Huang et al. (2022). Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process Cohen et al. (2019); Atzmon et al. (2020); Yan and Li (2024a, b). They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
2.2 Executing MAPF Plans
Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally. While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging Čáp et al. (2016), especially since MAPF plans are typically generated using imperfect robot models. Thus, we need techniques to execute MAPF plans safely given the potential delays. One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next. This ensures strict compliance with the MAPF plan but significantly reduces efficiency. Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well. Therefore, executing MAPF plans using an execution framework is preferred Čáp et al. (2016). The state-of-the-art execution framework is Action Dependency Graph (ADG) Hönig et al. (2019), which enforces passing orders defined by MAPF plans while accounting for delays. ADG can execute plans generated by MAPF planners of various simplified robot models Varambally et al. (2022).
An ADG, as shown in Fig. 2 (c), is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that encodes dependencies between actions from a MAPF plan. The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[1,n^{i}],r_{i}\in\mathcal{R}\}$ represents all actions performed by each robot $r_{i}$, where each vertex $v_{i}^{k}$ corresponds to the $k$-th action of $r_{i}$, and $n^{i}$ is the number of actions of $r_{i}$. The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which enforces the sequential execution of actions for each robot. For a robot $r_{i}$, a Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that action $v_{i}^{k}$ is completed before action $v_{i}^{k+1}$ starts. The edge set $\mathcal{E}_{2}$ includes Type-2 edges, which enforces the inter-robot passing order at shared vertices. For any two vertices $v_{i}^{k}$ and $v_{j}^{s}$ from $r_{i}$ and $r_{j}$, a Type-2 edge $(v_{i}^{k},v_{j}^{s})$ ensures that $v_{i}^{k}$ must be completed before $v_{j}^{s}$ can start. During execution, each action is performed according to these dependencies. The detailed execution procedure is described in Section 3.2.
2.3 Evaluating MAPF Plans
To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it. While many MAPF problem benchmark instances are proposed for discretized 2-D-grid maps Stern et al. (2019); Gebser et al. (2018); Sturtevant (2012); Mohanty et al. (2020); Qian et al. (2024), few software tools are provided to execute the MAPF plans in more realistic environments. Therefore, most prior works are evaluated using simplified robot models that assume no delays and discretized time and space.
Meanwhile, some software tools are presented to evaluate the MAPF plans in more realistic settings. MRP-Bench Schaefer et al. (2023) leverages a high-fidelity, physics-based simulator Gazebo Koenig and Howard (2004) to simulate multi-robot systems. It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays. REMROC Heuer et al. (2024) also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments. However, they have two main limitations. First, they often require a lot of engineering effort to use user-defined maps and MAPF planners, limiting their flexibility for different scenarios. Second, their scalability is constrained by the high computational overhead of Gazebo, which models many factors less relevant to MAPF research. MRP-Bench, for example, does not scale to more than 10 robots. Additionally, many MAPF researchers from the AI community may not be familiar with tools like Gazebo, making installation and usage a challenging process.
3 SMART Architecture
Fig. 3 shows an overview of SMART. It comprises three main modules: a simulator, an execution monitoring (EM) server, and executors. All modules are implemented in C++. The simulator generates simulation environments based on the user-provided configurations. The EM server parses the MAPF plan and monitors the execution status of each action throughout the simulation. When an action is ready for execution, the server sends the action to the corresponding robot via the network. The executor of each robot receives actions from the server, executes them using a controller, and synchronizes the execution status with the server accordingly.
3.1 Simulator
[…]
Figure 4: (a) Cycle in the ADG causing all robots to be stalled. (b) Adding safe stops between vertices to make the ADG acyclic.
3.2 Execution Monitoring (EM) Server
The EM server parses the MAPF plan and encodes its passing order into an ADG. It then manages the execution of the MAPF plan and communicates with the executors.
Server Initialization
During the initialization, we use the MAPF plan parser to construct an ADG. We pre-define the format to store the MAPF plan as a sequence of timed locations for each robot. The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans. This allows SMART to work with plan using varying simplified robot models. The parser first converts the path of each robot into a sequence of actions. Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path. If the required orientation differs from the current orientation, a rotate action is added. The ADG is constructed based on the sequence of actions. ADG requires the MAPF plan to be 1-robust, ensuring that robots do not collide even if delayed by one timestep. However, ensuring a 1-robust plan is not standard in most MAPF models. As a result, if we directly build an ADG based on a non-1-robust plan, we may create deadlocks, such as those illustrated in Fig. 4 (a). While these robots are planned to rotate simultaneously in a cycle in the non-1-robust plan, with ADG, they will end up waiting indefinitely for others to complete their next actions. To accommodate a broader range of planners, motivated by Hönig et al. (2016), we add an extra vertex between consecutive vertices in SMART as shown in Fig. 4 (b), making the ADG compatible with these MAPF models.
Execution Managing
During execution, an execution monitor framework manages the progress of execution and communicates with the executors. This framework follows the method described in ADG Hönig et al. (2019) to track and manage execution. Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished. At first, all vertices are staged. A vertex transitions to the enqueued if it has no preceding vertices or (1) its preceding vertex connected by a Type 1 edge is either enqueued or finished, and (2) its preceding vertices connected by Type 2 edges are finished. Actions associated with enqueued vertices are then sent to the executors for execution, while other vertices remain staged. A vertex is marked as finished when the executor confirms the completion of its associated action with the server. Communication with the executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
3.3 Executor
Each robot is launched with its own individual executor. As shown in Fig. 3, each executor consists of an RPC client, an action queue, and a controller. The RPC client establishes communication with the server, receiving enqueued actions and sending confirmation for finished actions. The action queue keeps track of the enqueued actions, which are executed in a first-in-first-out manner. A Proportional Integral Derivative (PID) controller is used to execute these actions by sending control commands to the simulator. Once an action is finished, a confirmation message is sent to the server. To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue.
3.4 Performance Metrics
SMART provides several built-in performance metrics, including (1) Average execution time (AET): The sum of the simulation times required by all robots to complete their assigned paths, divided by the number of robots, (2) Maximum execution time: The longest simulation time taken by any single robot to complete its path, and (3) Robot state over time: The state of an individual robot at different times, where the state of a robot includes the size of its action queue, its current position, and control output. Additionally, the open-source nature of our code allows researchers to easily derive custom metrics from SMART as needed. SMART also provides visualization for the MAPF plan executions, allowing researchers to better understand the behaviors of the robots.
4 System Evaluation
We test SMART on two machines: (1) a server with a 64-core AMD 7980X processor and 256 GB of memory, and (2) a laptop with a 4-core Intel i7-6700HQ processor and 16 GB of memory. Both machines are equipped with Ubuntu 20.04. We use machine (1) to evaluate the scalability and applicability of SMART in Sections 4.1 and 4.2. We use both machines to evaluate CPU and memory usage in Section 4.3.
4.1 Scalability and Runtime Performance
This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the total number of simulation updates to the real-world time (in seconds). A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands. In all experiments, the simulation update period is set to 0.1 simulation seconds. A higher simulation speed shows that more simulation updates can be finished within the same unit of time, indicating better scalability. For example, if the simulation speed is 100 and the update period is 0.1 simulation seconds, the simulation world runs 10 times faster than real time. We assess scalability on two environments converted from: warehouse (warehouse-10-20-10-2-1, size: 161$\times$63) from the MovingAI benchmark Stern et al. (2019) and sortation-center (size: $500\times 140$) from the League of Robot Runner MAPF competition Chan et al. (2024). MAPF plans are generated using MAPF-LNS2222Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 Li et al. (2022), a state-of-the-art MAPF planner, with a progressively increasing number of robots and a time limit of 60 seconds. These plans are then executed in SMART without visualization. Notably, the MAPF instances from warehouse support a maximum of 1,000 robots, whereas MAPF-LNS2 can handle up to 2,000 robots within the time limit on sortation-center.
Figure 5: Simulation speed on warehouse and sortation-center maps.
As shown in Fig. 5, SMART supports simulations with up to 2,000 robots. The more robots and the larger the map, the smaller the simulation speed. Nevertheless, SMART maintains a simulation speed of around 100 with 1,000 robots in warehouse. Even in the larger sortation-center, the simulation speed is consistently greater than 10.
Robots (#) average execution time maximum execution time
100 263.36 $\pm$ 0.01 469.80 $\pm$ 0.00
400 385.76 $\pm$ 0.02 737.60 $\pm$ 0.42
1000 612.04 $\pm$ 0.07 1389.43 $\pm$ 0.83
Table 1: Replicability using different numbers of robots.
4.2 Replicability
We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs. Due to the realistic setting we used, uncertainty may arise from factors such as network communication delays or controller variability. Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan. For each run, we record both the average execution time and the maximum execution time. As shown in Table 1, SMART demonstrates strong replicability across runs, with minor execution time variations. While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
4.3 CPU and Memory Usage
We evaluate the CPU and memory usage of SMART on both machines (1) and (2). We use machine (2) to demonstrate that SMART can run on resource-constrained machines. To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark Stern et al. (2019). All simulations are parallelized on all CPU cores on both machines. We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots. Table 2 shows the results. We observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops. In machine (2), the simulation speed is slower than in machine (1) due to an inferior processor. For reference, we run the Gazebo-based simulator Varambally et al. (2022) on machine (2) with 50 robots in an environment of size 31$\times$12. Since Gazebo operates on a single thread, it utilizes 100% of a single CPU core, consumes 2.09 GB of memory, and achieves a simulation speed of 5.40.
Machine Robots (#) CPU Usage (%) Memory Usage (GB) Simulation Speed
(1) 100 16.48 $\pm$ 0.58 0.19 $\pm$ 0.26 334.90 $\pm$ 1.54
1000 46.02 $\pm$ 0.98 0.65 $\pm$ 0.03 73.71 $\pm$ 0.40
(2) 100 27.96 $\pm$ 0.39 0.04 $\pm$ 0.00 143.74 $\pm$ 1.03
1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 19.45 $\pm$ 0.09
Table 2: Comparison of CPU and Memory Usage
5 Problems from MAPF to Realistic Scenarios
We argue that SMART has several broader impacts on the MAPF research community. First, it allows researchers to evaluate MAPF plans from arbitrary MAPF planners in a simulator with realistic robot kinodynamics models and execution delays caused by real-world factors. Second, it bridges the gap between MAPF planning algorithms and executing the planned paths in the real world, bringing researchers more guidance on designing high-quality MAPF algorithms for the real world. Third, SMART can be used to study research problems regarding executing MAPF plans in the real world.
In this section, we study research problems focusing on the gap between MAPF planning and real-world execution of the planned paths. We study these problems using six simulation environments converted from six maps from the MovingAI benchmark Stern et al. (2019): empty (empty-32-32, size: 32$\times$32), random (random-64-64-10, size: 64$\times$64), room (room-64-64-16, size: 64$\times$64), den312d (den312d, size: 65$\times$81), maze (maze-32-32-4, size: 32$\times$32), and warehouse (warehouse-10-20-10-2-1, size: 161$\times$63).
Feature All Features SOC Rotations (#) Type-1 edges (#) Type-2 edges (#) Conflict robot pairs (#) Robots (#)
MAPE 0.0342 0.1778 0.1793 0.2414 0.2826 0.3175 0.4317
Table 3: Quadratic Regression Results. Except for the first column, each column means that the model is trained with only a single feature
5.1 Impact of Optimality
We first use SMART to study the practical effectiveness of the objective used by MAPF algorithms. One of the most widely used objectives in MAPF research is the Sum of Costs (SOC), which measures the total number of timesteps for all robots to reach their goals. While extensive efforts have been made to develop algorithms that optimize SOC optimally or sub-optimally, it is unclear, for example, whether reducing SOC by 1% translates to a similar degree of improvement in AET in a more realistic simulation. In this section, we use SMART to assess the correlation between improvements in SOC and AET. Moreover, we want to explore better objectives than SOC in terms of evaluating the quality of the MAPF plan in more realistic settings.
5.1.1 Experiment Setup
We generate MAPF plans with varying SOC for a specific MAPF instance by using MAPF-LNS [^3] Code at https://github.com/Jiaoyang-Li/MAPF-LNS Li et al. (2021a), an anytime method that iteratively refines MAPF plans to improve SOC within a predefined runtime limit. For each map, we select three different numbers of robots and five sets of different start and goal locations for robots. For each MAPF instance, we record all plans generated in each iteration of MAPF-LNS within the runtime limit of 60 seconds. We then execute these plans in SMART.
Figure 6: Relationship between SOC and AET. Each dot represents a MAPF plan and its execution result.
5.1.2 Results and Analysis
We first study whether SOC can measure the real-world execution time of a MAPF plan. As shown in Fig. 6, SOC and AET show a strong positive correlation across all environments and all numbers of robots, indicating that SOC captures the trend of AET. However, the correlation is not perfectly monotonic. For the same SOC value, there are noticeable variations in AET, suggesting that SOC alone cannot account for all factors influencing execution time. This highlights the limitation of SOC as the sole objective of the MAPF problem.
To explore other factors that influence execution time, we extract additional features from the ADG, including (1) Type-2 edges (#), the number of type-2 edges, (2) Type-1 edges (#), the number of type-1 edges, (3) Rotations (#), the number of rotation actions, (4) Conflict robot pairs (#), the number of robot pairs connected by type-2 edges, reflecting dependencies in their passing orders, and (5) Robots (#), the number of robots. We analyze the correlation between these features and AET by using a quadratic regression model to predict AET based on these features. We measure prediction accuracy by using the mean absolute percentage error (MAPE). As shown in Table 3, SOC exhibits the highest predictive power among individual features, while other features, such as the number of Type-2 edges, Type-1 edges, and Rotations, also show correlation with AET. Meanwhile, the model incorporating all features achieves a much lower MAPE than any single feature, indicating that multiple features influence the execution time. In conclusion, incorporating features in addition to SOC is crucial to developing a more accurate objective function for execution time estimation during MAPF planning.
Figure 7: AET comparison for MAPF with different robot models.
5.2 Impact of Different Robot Models
The standard MAPF problem relies on a simplified model that ignores critical execution factors such as kinodynamics and potential delays. This simplification creates a gap between planning and real-world execution. Thus, recent research has introduced more realistic MAPF models that consider real-world factors. An ideal MAPF model should capture the essential aspects of robot kinodynamics while omitting unnecessary details. As SMART supports various robot models, we conduct an empirical comparison of different MAPF models to identify the most effective one in real-world scenarios. In Varambally et al. (2022), a similar problem is studied, but the experiments are limited to a warehouse map and consider no more than 50 robots. Additionally, we compare a wider range of robot models.
5.2.1 Experiment Setup
We evaluate the impact of incorporating robot kinodynamics and potential delays. For the kinodynamics models, we consider: (1) the standard MAPF model, (2) the MAPF model with rotation, and (3) the MAPF model that accounts for the full kinodynamics of robots, including rotation, speed, and acceleration. To model delays, we use the $k$-robust delay model Atzmon et al. (2020), which ensures that robots can avoid collisions even if delayed by up to $k$ timesteps. By combining kinodynamics and delay factors, we evaluate the following MAPF algorithms: (1) MAPF, $k=i$: This method uses the standard MAPF model and a $k$-robust delay model of $k=i$. (2) MAPF+rotation, $k=i$: This extends the standard MAPF model by incorporating in-place rotation but does not account for other kinematics. It also includes a $k$-robust delay model. (3) MAPF+kinodynamics: Derived from Yan and Li (2024a), this method incorporates accurate kinodynamic constraints of robots. We use PBS Ma et al. (2019) to handle conflicts between robots under these models with a runtime limit of 60 seconds. We run these methods to generate MAPF plans for all maps using 25 “random scenarios” with a progressive number of robots until the success rate is smaller than 20%. Finally, these plans are executed using SMART.
Figure 8: Paths generated on different maps.
5.2.2 Results and Analysis
As shown in Fig. 7, the method incorporating the full kinodynamics of robots achieves the shortest AET among all the models. Additionally, the planner considering robot rotations demonstrates much better AET than the standard MAPF model. These results indicate that incorporating more accurate kinodynamic models into planning can improve execution time. However, we also observe trade-offs associated with these models. More complex kinodynamic models lead to longer runtimes during planning. Meanwhile, the $k$-robust model achieves better or comparable performance with fewer robots with larger values of $k$; however, its performance declines significantly as the number of robots increases, resulting in lower solution quality and reduced scalability. Interestingly, in warehouse environments, MAPF+kinodynamics does not provide significant improvements over the MAPF+rotation. To better understand this, we visualize the MAPF plans in Fig. 8, comparing the warehouse map with random and room maps. The planned paths show that MAPF+rotation and MAPF+kinodynamics generate paths with more consecutive movements and fewer conflict vertices in warehouse maps than in the other two maps. We hypothesize that this minimizes the potential benefits of accounting for kinodynamics in such environments. In summary, while more accurate kinodynamic models generally improve execution time in realistic settings, their increased computational complexity can limit scalability. Additionally, the $k$-robust model is less efficient and may even harm performance, particularly in more congested scenarios.
Figure 9: AET using different MAPF models and planners.
5.3 Optimality and Robot Model
From Sections 5.1 and 5.2, we show that both improving SOC and incorporating a more accurate kinodynamic model during planning can lead to better AET during execution. However, both of them require additional computation time. Considering the limited planning time available in real-world applications, this section explores the balance between these factors to provide practical insights.
5.3.1 Experiment Setup
We generate MAPF plans with varying SOC for three robot models: (1) standard MAPF, (2) MAPF with rotation, and (3) MAPF with kinodynamics. To obtain optimal solutions, we use CBS [^4] Code at https://github.com/Jiaoyang-Li/CBSH2-RTC Li et al. (2021b) for the standard MAPF model and CBS with rotation [^5] Code at https://github.com/YueZhang-studyuse/MAPF_T Zhang et al. (2023a) for the rotation model. For suboptimal plans, we combine PBS with all three models, as described in Section 5.2. To analyze the relationship between AET and SOC, we create a naive anytime planner by running PP with random restarts and recording the MAPF plan with the lowest SOC until reaching a runtime limit of 60 seconds. We do not use existing anytime solvers, such as MAPF-LNS, because, to our best knowledge, no available implementation considers robot models with kinodynamic constraints.
5.3.2 Results and Analysis
As shown in Fig. 9, planners using more accurate robot models consistently achieve better AET across all maps, even when the methods using less accurate models are optimal. This indicates that incorporating more accurate kinodynamic robot models into planning may be more important than obtaining optimal solutions with less accurate models. Notably, as discussed in Section 5.1, since SOC is not a perfect metric, the optimal planner does not always achieve the best AET. Meanwhile, execution frameworks may also become a bottleneck. Since ADG adopts an overly conservative strategy, assuming robots can be delayed indefinitely, it compromises the solution quality of the MAPF plan in certain cases. In conclusion, the limiting factors for achieving better AET lie in the more accurate robot models or the execution frameworks with better performance.
6 Conclusion
We introduced SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings. SMART bridges the gap between methods that use simplified models and their real-world deployments. SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments. Our experiments demonstrated that SMART can handle thousands of robots while maintaining high simulation speeds. We highlight several key challenges arising from the application of MAPF in realistic scenarios. We further show how SMART can be used to address these challenges. Future work includes extending the framework to address lifelong MAPF scenarios, integrating it with other execution frameworks Coskun and O’Kane (2019); Berndt et al. (2023); Feng et al. (2024), and interfacing it with real robots.
Acknowledgments
The research was supported by the National Science Foundation (NSF) under grant number #2328671 and a gift from Amazon. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.

### v2

@@
 Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings.
-SMART fills this gap with several advantages: (1) SMART uses a physics-engine-based simulator to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots.
-In addition, we use SMART to explore and demonstrate research questions about the execution of MAPF algorithms in real-world scenarios.
-The code is publicly available at https://jingtianyan.github.io/publication/2025-smart.
-1 Introduction
-Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c) Robot model used in the environment. (d) A 3D view of the environment.
-The Multi-Agent Path Finding (MAPF) problem Stern et al.
-(2019) focuses on planning collision-free paths for large groups of robots within a known environment.
-State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds Ma et al.
-(2019); Li et al.
-(2021c); Zhang et al.
-(2022); Li et al.
-(2022).
+SMART fills this gap with several advantages: (1) SMART uses physics-engine-based simulators to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots.
+The code is publicly available at https://github.com/smart-mapf/smart.
+Index Terms:
+Multi-Agent Path Finding, Multi-Robot System, Realistic Simulation.
+I Introduction
+The Multi-Agent Path Finding (MAPF) problem [1] focuses on planning collision-free paths for a large group of robots within a known environment.
+This problem finds various real-world applications [2, 3, 4], with warehouse automation being a prominent example, where hundreds of robots must be coordinated to perform daily operations.
+In recent years, MAPF has grown into a rapidly expanding research area [5, 6].
+State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds [7, 8, 9].
 However, these algorithms make several simplifying assumptions.
-First, they often rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths.
+First, they rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths.
 Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors.
-Third, they typically plan and evaluate paths on unrealistic discretized 2-D grid maps Stern et al.
-(2019).
-Therefore, while researchers have shown that MAPF algorithms have a wide range of applications Ma et al.
-(2017); Hönig et al.
-(2019); Ho et al.
-(2022); Li et al.
-(2023); Zhang et al.
-(2023b, c, 2024); Zang et al.
-(2025), the algorithms are mostly evaluated under simplified settings, leaving their performance in more realistic settings unclear.
+Therefore, while researchers have shown that MAPF algorithms have a wide range of applications, their performance in more realistic settings is unclear.
+Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c, d) Robot models used in the environment. (e) A 3D view of the environment.
 MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings.
@@
 Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort.
-While some software tools Schaefer et al.
-(2023); Heuer et al.
-(2024) have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps.
+While some software tools [10, 11] have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps.
 In addition, industrial professionals possess realistic simulators for their applications of interest.
@@
 In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios.
-SMART comprises three main components: (1) a physics-engine-based simulator based on Argos 3 Pinciroli et al.
-(2012), (2) an execution monitoring (EM) server, and (3) robot-specific executors.
+SMART comprises three main components: (1) physics-engine-based simulators, (2) an execution monitoring (EM) server, and (3) robot-specific executors.
 The simulator generates realistic simulation environments based on user-defined configurations.
 Fig.
-1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark Stern et al.
-(2019).
+1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark [1].
 The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors.
-It leverages the Action Dependency Graph (ADG) Hönig et al.
-(2019), a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models.
+It leverages the Action Dependency Graph (ADG) [3], a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models.
 Each robot is assigned an individual executor that receives and executes actions from the server.
-SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and spatial inaccuracies caused by controller limitations.
-Our contributions in this paper are: (1) We present SMART, an open-source software tool to evaluate MAPF algorithms in realistic scenarios, which can scale to thousands of robots while ensuring replicability, (2) we evaluate SMART with a variety of open-source implementations of MAPF algorithms and robot models, showing its flexibility, and (3) we show some important findings regarding the application of MAPF algorithms in realistic scenarios.
-We also use them as examples of how SMART can be used to address such problems.
+SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and temporal uncertainty during execution.
+Our contributions in this paper are as follows: 1.
+We present SMART, an open-source software tool for evaluating MAPF algorithms in realistic scenarios.
+SMART directly accepts MAPF plans and problem instances from existing benchmarks.
+2.
+We integrate SMART with two simulators, ARGoS3 and Isaac Sim, and robot models and empirically evaluate its scalability and replicability.
+SMART can scale to thousands of robots while maintaining consistent and reproducible results, while prior tools fail to handle a hundred robots.
+3.
+We develop an online interface with interactive visualization, allowing users to easily monitor execution progress and access detailed statistical data during simulation.
+4.
+We validate SMART on real-world mobile robots, demonstrating its robustness and effectiveness in physical environments.
 Figure 2: (a) MAPF problem.
@@
 Figure 3: System overview of SMART.
-2 Background
-2.1 MAPF Problem and MAPF Plan
-A MAPF instance Stern et al.
-(2019) consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ and a set of $I$ robots =$\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations.
+II Background
+In this section, we first introduce the MAPF problem and related work.
+Next, we discuss existing execution frameworks for executing MAPF plans.
+Finally, we review benchmarks used to evaluate MAPF algorithms.
+II-A MAPF Problem and MAPF Plan
+A MAPF instance [1] consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ constructed from a 2-D grid map and a set of $I$ robots $\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations.
 Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations.
@@
 Two robots collide if they are at the same vertex or swap vertices at the same timestep.
-An example MAPF instance and plan on a 2-D grid graph is shown in Fig.
+An example MAPF instance and plan on a 2-D grid graph are shown in Fig.
 2 (a) and (b), respectively.
 The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
-Prior research has proposed many MAPF algorithms.
-Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees Sharon et al.
-(2015); Barer et al.
-(2014); Li et al.
-(2021c), while others focus on scalability, solving MAPF problems involving thousands of robots Okumura et al.
-(2022); Okumura (2023).
-Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency Li et al.
-(2021a); Huang et al.
-(2022).
-Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process Cohen et al.
-(2019); Atzmon et al.
-(2020); Yan and Li (2024a, b).
+Prior research has proposed many MAPF algorithms, most of which are developed based on 2-D grid maps.
+Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees [12, 13], while others focus on scalability, solving MAPF problems involving thousands of robots [14, 15].
+Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency [16, 17].
+Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process [18, 19, 20, 21].
 They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
-2.2 Executing MAPF Plans
+II-B Executing MAPF Plans
 Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally.
-While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging Čáp et al.
-(2016), especially since MAPF plans are typically generated using imperfect robot models.
-Thus, we need techniques to execute MAPF plans safely given the potential delays.
+While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging [22], especially since MAPF plans are typically generated using imperfect robot models.
+Thus, we need techniques to execute MAPF plans safely given the uncertainty in execution time.
 One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next.
@@
 Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well.
-Therefore, executing MAPF plans using an execution framework is preferred Čáp et al.
-(2016).
-The state-of-the-art execution framework is Action Dependency Graph (ADG) Hönig et al.
-(2019), which enforces passing orders defined by MAPF plans while accounting for delays.
-ADG can execute plans generated by MAPF planners of various simplified robot models Varambally et al.
-(2022).
+Therefore, executing MAPF plans using an execution framework is preferred [22].
+The state-of-the-art execution framework is Action Dependency Graph (ADG) [3], which enforces passing orders defined by MAPF plans while accounting for delays.
+ADG can execute plans generated by MAPF planners of various robot models [23].
 An ADG, as shown in Fig.
 2 (c), is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that encodes dependencies between actions from a MAPF plan.
-The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[1,n^{i}],r_{i}\in\mathcal{R}\}$ represents all actions performed by each robot $r_{i}$, where each vertex $v_{i}^{k}$ corresponds to the $k$-th action of $r_{i}$, and $n^{i}$ is the number of actions of $r_{i}$.
-The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which enforces the sequential execution of actions for each robot.
+The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[1,n^{i}],r_{i}\in\mathcal{R}\}$ represents all actions performed by each robot $r_{i}$, where each vertex $v_{i}^{k}$ corresponds to the $k$-th action of $r_{i}$, and $n^{i}$ is its number of actions.
+The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which enforce the sequential execution of actions by each robot.
 For a robot $r_{i}$, a Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that action $v_{i}^{k}$ is completed before action $v_{i}^{k+1}$ starts.
-The edge set $\mathcal{E}_{2}$ includes Type-2 edges, which enforces the inter-robot passing order at shared vertices.
+The edge set $\mathcal{E}_{2}$ includes Type-2 edges, which enforce the inter-robot passing order at shared vertices.
 For any two vertices $v_{i}^{k}$ and $v_{j}^{s}$ from $r_{i}$ and $r_{j}$, a Type-2 edge $(v_{i}^{k},v_{j}^{s})$ ensures that $v_{i}^{k}$ must be completed before $v_{j}^{s}$ can start.
 During execution, each action is performed according to these dependencies.
-The detailed execution procedure is described in Section 3.2.
-2.3 Evaluating MAPF Plans
+The detailed execution procedure is described in Section III-B.
+II-C Evaluating MAPF Plans
 To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it.
-While many MAPF problem benchmark instances are proposed for discretized 2-D-grid maps Stern et al.
-(2019); Gebser et al.
-(2018); Sturtevant (2012); Mohanty et al.
-(2020); Qian et al.
-(2024), few software tools are provided to execute the MAPF plans in more realistic environments.
-Therefore, most prior works are evaluated using simplified robot models that assume no delays and discretized time and space.
-Meanwhile, some software tools are presented to evaluate the MAPF plans in more realistic settings.
-MRP-Bench Schaefer et al.
-(2023) leverages a high-fidelity, physics-based simulator Gazebo Koenig and Howard (2004) to simulate multi-robot systems.
+While many MAPF benchmarks with 2-D-grid maps are proposed [1, 24, 25, 26], most prior MAPF works are evaluated on them using simplified robot models that assume no delays and discretized time and space.
+Some benchmarks have attempted to include realistic elements such as acceleration and rotation [27, 28].
+However, their model is still far from realistic, and typically requires the same robot models for both planning and execution phases.
+Given the difficulty of modeling all real-world factors analytically, physics-engine-based simulators offer a promising approach for more accurate evaluation.
+MRP-Bench [10] leverages a high-fidelity, physics-based simulator Gazebo [29] to simulate multi-robot systems.
 It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays.
-REMROC Heuer et al.
-(2024) also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments.
-However, they have two main limitations.
-First, they often require a lot of engineering effort to use user-defined maps and MAPF planners, limiting their flexibility for different scenarios.
-Second, their scalability is constrained by the high computational overhead of Gazebo, which models many factors less relevant to MAPF research.
+REMROC [11] also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments.
+However, these existing tools have two main limitations.
+First, they often require significant engineering effort to integrate user-defined maps and MAPF planners, limiting their flexibility for different scenarios.
+In particular, they lack a general execution framework, making it difficult to directly evaluate MAPF planners using different robot models.
+Second, they have limited scalability: the largest number of robots tested in previous work is less than 100 agents.
+Most of these works rely on Gazebo, whose high computational overhead, resulting from modeling many factors not relevant to MAPF, significantly constrains scalability.
 MRP-Bench, for example, does not scale to more than 10 robots.
-Additionally, many MAPF researchers from the AI community may not be familiar with tools like Gazebo, making installation and usage a challenging process.
-3 SMART Architecture
+III SMART Architecture
 Fig.
@@
 The executor of each robot receives actions from the server, executes them using a controller, and synchronizes the execution status with the server accordingly.
-3.1 Simulator
-[…]
+III-A Simulator
+Given a user-provided configuration, the simulator creates realistic simulation environments.
+The configuration specifies obstacle locations and shapes, robot start locations, and detailed robot configurations (e.g., the robot model, as well as its velocity and acceleration limits).
+To support the widely used MovingAI benchmark [1], SMART can converts its 2-D grid maps and corresponding MAPF instances into the configuration required by the simulator.
+During execution, the simulator uses a physics engine to update the states of the robots and environment, including their poses, dynamics, sensors, and actuators, based on the received control commands from the executors.
+The executors can query the simulator for real-time robot information, such as their current pose.
+To accommodate diverse user requirements, SMART supports two simulation platforms: ARGoS3 [30] and Isaac Sim [31].
+III-A1 ARGoS3
+ARGoS3 is a multi-robot simulator with excellent scalability, advanced physics engines, and support for custom robot and environment models, making it suitable for our tasks.
+While users can specify any robot model, we use the foot-bot (shown in Fig.
+1 (c)), a differential-drive robot, in our experiments.
+It closely resembles robots used in real-world warehouses and can move forward and rotate in place.
+III-A2 Isaac Sim
+Isaac Sim is a high-fidelity robotics simulator developed by NVIDIA, featuring GPU-accelerated physics, photorealistic rendering, and seamless integration with ROS.
+It is well-suited for sim-to-real transfer and physically grounded perception and control tasks.
+We use the JetBot robot (shown in Fig.
+1 (d)), a differential-drive robot commonly used in research and education.
 Figure 4: (a) Cycle in the ADG causing all robots to be stalled. (b) Adding safe stops between vertices to make the ADG acyclic.
-3.2 Execution Monitoring (EM) Server
+III-B Execution Monitoring (EM) Server
 The EM server parses the MAPF plan and encodes its passing order into an ADG. It then manages the execution of the MAPF plan and communicates with the executors.
@@
 During the initialization, we use the MAPF plan parser to construct an ADG.
-We pre-define the format to store the MAPF plan as a sequence of timed locations for each robot.
+We define a standardized format to store the MAPF plan as a sequence of timed locations for each robot $r_{i}$, i.e., $((x_{i}^{0},y_{i}^{0},t_{i}^{0}),...,(x_{i}^{T},y_{i}^{T},t_{i}^{T}))$.
 The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans.
-This allows SMART to work with plan using varying simplified robot models.
+This allows SMART to work with plans using varying simplified robot models.
 The parser first converts the path of each robot into a sequence of actions.
-Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path.
-If the required orientation differs from the current orientation, a rotate action is added.
+Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path and adds a rotation action when needed.
 The ADG is constructed based on the sequence of actions.
-ADG requires the MAPF plan to be 1-robust, ensuring that robots do not collide even if delayed by one timestep.
-However, ensuring a 1-robust plan is not standard in most MAPF models.
-As a result, if we directly build an ADG based on a non-1-robust plan, we may create deadlocks, such as those illustrated in Fig.
+ADG requires the MAPF plan to be circle-conflict free, ensuring that robots do not collide even if delayed by one timestep [3].
+However, ensuring this is not standard in most MAPF models.
+As a result, if we directly build an ADG based on a plan with circle conflict, we may create deadlocks, such as the one illustrated in Fig.
 4 (a).
-While these robots are planned to rotate simultaneously in a cycle in the non-1-robust plan, with ADG, they will end up waiting indefinitely for others to complete their next actions.
-To accommodate a broader range of planners, motivated by Hönig et al.
-(2016), we add an extra vertex between consecutive vertices in SMART as shown in Fig.
+While these robots are planned to rotate simultaneously in a cycle, with ADG, they will end up waiting indefinitely for others to complete their next actions.
+To accommodate a broader range of planners, motivated by [32], we add an extra vertex between consecutive vertices in SMART as shown in Fig.
 4 (b), making the ADG compatible with these MAPF models.
 Execution Managing
-During execution, an execution monitor framework manages the progress of execution and communicates with the executors.
-This framework follows the method described in ADG Hönig et al.
-(2019) to track and manage execution.
+During execution, an execution monitor framework manages the progress of execution and communicates with robot executors.
+This framework follows the method described in ADG [3] to track and manage execution.
 Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished.
 At first, all vertices are staged.
-A vertex transitions to the enqueued if it has no preceding vertices or (1) its preceding vertex connected by a Type 1 edge is either enqueued or finished, and (2) its preceding vertices connected by Type 2 edges are finished.
-Actions associated with enqueued vertices are then sent to the executors for execution, while other vertices remain staged.
+A vertex transitions to enqueued if (1) it has no preceding vertices, or (2) its preceding vertex connected by a Type-1 edge is enqueued or finished, and all its preceding vertices connected by Type-2 edges are finished.
+Actions associated with enqueued vertices are then sent to the executors for execution.
 A vertex is marked as finished when the executor confirms the completion of its associated action with the server.
-Communication with the executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
-3.3 Executor
-Each robot is launched with its own individual executor.
+Communication with robot executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
+III-C Executor
+Each robot is launched with an individual executor.
 As shown in Fig.
@@
 To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue.
-3.4 Performance Metrics
-SMART provides several built-in performance metrics, including (1) Average execution time (AET): The sum of the simulation times required by all robots to complete their assigned paths, divided by the number of robots, (2) Maximum execution time: The longest simulation time taken by any single robot to complete its path, and (3) Robot state over time: The state of an individual robot at different times, where the state of a robot includes the size of its action queue, its current position, and control output.
+For example, if a robot receives three consecutive forward movement actions, the executor merges them into a single longer forward action, reducing unnecessary stop-and-go behavior.
+III-D Performance Metrics
+SMART provides several built-in performance metrics, including (1) Average execution time (AET): The sum of the simulation times required by all robots to complete their assigned paths, divided by the number of robots, (2) Maximum execution time: The longest simulation time taken by any single robot to complete its path, and (3) Robot state over time: The state of each robot at different times, where the state includes the size of its action queue, its current position, and its current wheel velocity.
 Additionally, the open-source nature of our code allows researchers to easily derive custom metrics from SMART as needed.
 SMART also provides visualization for the MAPF plan executions, allowing researchers to better understand the behaviors of the robots.
-4 System Evaluation
-We test SMART on two machines: (1) a server with a 64-core AMD 7980X processor and 256 GB of memory, and (2) a laptop with a 4-core Intel i7-6700HQ processor and 16 GB of memory.
+IV Using SMART
+Users can evaluate their solutions using SMART in two ways: by uploading their MAPF problem and solution through a web interface or by deploying SMART locally.
+Figure 5: The SMART web interface visualizing a multi-robot simulation on a grid map.
+In the viewport, users interact with individual robots to inspect paths and inter-robot constraints.
+On the side panel, users can configure the environment, control playback, and see execution details.
+SMART as a Service
+SMART, with ARGoS3 and the 3D visualizer, is directly available as software-as-a-service [^2] The link will be made available here upon acceptance..
+Specifically, we provide an intuitive web app for running and inspecting MAPF simulations.
+Complete with a beautiful 3D-visualizer built with React and Three.js, it allows researchers to instantly understand how algorithms perform with physical robots (Fig.
+5).
+Users configure the environment by specifying a map and scenario file in the MovingAI Benchmark format.
+Within the online interface, after users upload the files and select the desired kinodynamic settings for the robots, and initiate the simulation with a single click.
+After starting the simulation, they can control the simulation time step and adjust the playback speed.
+Users may select robots to view complete paths, execution status, and Type-2 edges between robots.
+Installing SMART Locally
+Since SMART provides flexible interfaces for customization, researchers may choose to compile it locally for advanced use cases, such as running it in headless mode, integrating it into custom pipelines, or interfacing with real robots.
+As an open-source tool, SMART includes documentation and examples to support these advanced configurations.
+V System Evaluation
+We test SMART on two machines.
+All experiments are conducted on machine (1), a high-performance server with a 64-core AMD 7980X processor, 256 GB of memory, and RTX 3090Ti GPU.
+In Section V-C, we additionally test SMART on machine (2), a low-end laptop with a 4-core Intel i7-6700HQ processor and 16 GB of memory, to demonstrate SMART’s performance on resource-constrained hardware.
 Both machines are equipped with Ubuntu 20.04.
-We use machine (1) to evaluate the scalability and applicability of SMART in Sections 4.1 and 4.2.
-We use both machines to evaluate CPU and memory usage in Section 4.3.
-4.1 Scalability and Runtime Performance
-This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the total number of simulation updates to the real-world time (in seconds).
+Figure 6: Simulation environments. (a) empty (b) maze (c)random (d) game (e) warehouse (f) sortation-center.
+[…]
+In all experiments, the simulation update period is set to 0.1 simulation seconds for both simulators.
 A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands.
-In all experiments, the simulation update period is set to 0.1 simulation seconds.
-A higher simulation speed shows that more simulation updates can be finished within the same unit of time, indicating better scalability.
-For example, if the simulation speed is 100 and the update period is 0.1 simulation seconds, the simulation world runs 10 times faster than real time.
-We assess scalability on two environments converted from: warehouse (warehouse-10-20-10-2-1, size: 161$\times$63) from the MovingAI benchmark Stern et al.
-(2019) and sortation-center (size: $500\times 140$) from the League of Robot Runner MAPF competition Chan et al.
-(2024).
-MAPF plans are generated using MAPF-LNS2222Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 Li et al.
-(2022), a state-of-the-art MAPF planner, with a progressively increasing number of robots and a time limit of 60 seconds.
-These plans are then executed in SMART without visualization.
-Notably, the MAPF instances from warehouse support a maximum of 1,000 robots, whereas MAPF-LNS2 can handle up to 2,000 robots within the time limit on sortation-center.
-Figure 5: Simulation speed on warehouse and sortation-center maps.
+In ARGoS3, each grid cell represents 1 meter.
+Each foot-bot robot is subject to a speed limit of [0, 5] m/s, an acceleration limit of [-0.4, 0.4] m/s2, and an angular velocity limit of 30/s.
+In Isaac Sim, each grid cell represents 0.5 meters.
+Each JetBot is constrained to a speed limit of [0, 2.5] m/s, an acceleration limit of [-0.2, 0.2] m/s2, and the same angular velocity limit of 30/s.
+We evaluated our system using MAPF plans generated by three different methods: CBS [12], PBS [7], and MAPF-LNS2 [9].
+These plans were created using four different robot models: the standard MAPF model [1], the $k$-robust delay model [19], MAPF with rotation [33], and MAPF with kinodynamics [21].
+Across all combinations, our system achieved a 100% execution success rate.
+In the remainder of this section, we present results based on MAPF plans generated by MAPF-LNS2, as it demonstrated the highest scalability among the algorithms evaluated.
+V-A Scalability and Runtime Performance
+This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the simulation time to the real-world time.
+A higher simulation speed means that more simulation updates can be finished within the same unit of time, indicating better scalability.
+For example, if the simulation speed is 10, the simulation world runs 10 times faster than in real time.
+MAPF plans are generated using MAPF-LNS2333Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 [9] with a progressively increasing number of robots and a time limit of 60 seconds.
+For each map, we randomly select 5 scenarios from the MovingAI Benchmark to generate plans, which are then executed in SMART without visualization.
+In ARGoS3, the maximum number of robots tested corresponds to the limit imposed by the MAPF instances from the MovingAI Benchmark for all maps except sortation-center, which is set to 2,000 robots.
+In contrast, for Isaac Sim, the upper bound on the number of robots is determined by the point at which the simulation speed drops below real-time (i.e., simulation speed << 1).
+Simulator Robots average execution maximum execution
+(#) time time
+ARGoS3 10 123.33 $\pm$ 0.00 373.6 $\pm$ 0.00
+50 182.84 $\pm$ 0.01 433.30 $\pm$ 0.00
+100 263.36 $\pm$ 0.01 469.80 $\pm$ 0.00
+200 269.79 $\pm$ 0.02 675.50 $\pm$ 0.00
+400 385.76 $\pm$ 0.02 737.60 $\pm$ 0.42
+600 462.87 $\pm$ 0.07 843.70 $\pm$ 0.71
+1000 612.04 $\pm$ 0.07 1389.43 $\pm$ 0.83
+Isaac Sim 10 122.47 $\pm$ 0.00 369.47 $\pm$ 0.00
+50 183.71 $\pm$ 0.07 435.46 $\pm$ 1.36
+TABLE I: Replicability using different numbers of robots.
+Figure 7: Simulation speed on all environments.
 As shown in Fig.
-5, SMART supports simulations with up to 2,000 robots.
+7, SMART supports simulations with up to 2,000 robots.
 The more robots and the larger the map, the smaller the simulation speed.
-Nevertheless, SMART maintains a simulation speed of around 100 with 1,000 robots in warehouse.
-Even in the larger sortation-center, the simulation speed is consistently greater than 10.
-Robots (#) average execution time maximum execution time
-100 263.36 $\pm$ 0.01 469.80 $\pm$ 0.00
-400 385.76 $\pm$ 0.02 737.60 $\pm$ 0.42
-1000 612.04 $\pm$ 0.07 1389.43 $\pm$ 0.83
-Table 1: Replicability using different numbers of robots.
-4.2 Replicability
+Nevertheless, for ARGoS3, SMART maintains a simulation speed of around 10.0 with 1,000 robots in warehouse.
+Even in the larger sortation-center, the simulation speed is consistently greater than 1.0.
+In contrast, Isaac Sim provides the capability to model more complex real-world factors while also producing photorealistic visualizations, but it is significantly less scalable, particularly in large environments.
+It is therefore more suitable for experiments with a smaller number of agents.
+V-B Replicability
 We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs.
@@
 Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan.
-For each run, we record both the average execution time and the maximum execution time.
-As shown in Table 1, SMART demonstrates strong replicability across runs, with minor execution time variations.
+For each run, we record both the average and the maximum execution time.
+As shown in Table I, SMART demonstrates strong replicability across runs, with minor execution time variations.
 While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
-4.3 CPU and Memory Usage
+V-C CPU and Memory Usage
 We evaluate the CPU and memory usage of SMART on both machines (1) and (2).
 We use machine (2) to demonstrate that SMART can run on resource-constrained machines.
-To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark Stern et al.
-(2019).
+To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark [1].
 All simulations are parallelized on all CPU cores on both machines.
 We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots.
-Table 2 shows the results.
-We observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops.
-In machine (2), the simulation speed is slower than in machine (1) due to an inferior processor.
-For reference, we run the Gazebo-based simulator Varambally et al.
-(2022) on machine (2) with 50 robots in an environment of size 31$\times$12.
-Since Gazebo operates on a single thread, it utilizes 100% of a single CPU core, consumes 2.09 GB of memory, and achieves a simulation speed of 5.40.
-Machine Robots (#) CPU Usage (%) Memory Usage (GB) Simulation Speed
-(1) 100 16.48 $\pm$ 0.58 0.19 $\pm$ 0.26 334.90 $\pm$ 1.54
-1000 46.02 $\pm$ 0.98 0.65 $\pm$ 0.03 73.71 $\pm$ 0.40
-(2) 100 27.96 $\pm$ 0.39 0.04 $\pm$ 0.00 143.74 $\pm$ 1.03
-1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 19.45 $\pm$ 0.09
-Table 2: Comparison of CPU and Memory Usage
-5 Problems from MAPF to Realistic Scenarios
-We argue that SMART has several broader impacts on the MAPF research community.
-First, it allows researchers to evaluate MAPF plans from arbitrary MAPF planners in a simulator with realistic robot kinodynamics models and execution delays caused by real-world factors.
-Second, it bridges the gap between MAPF planning algorithms and executing the planned paths in the real world, bringing researchers more guidance on designing high-quality MAPF algorithms for the real world.
-Third, SMART can be used to study research problems regarding executing MAPF plans in the real world.
-In this section, we study research problems focusing on the gap between MAPF planning and real-world execution of the planned paths.
-We study these problems using six simulation environments converted from six maps from the MovingAI benchmark Stern et al.
-(2019): empty (empty-32-32, size: 32$\times$32), random (random-64-64-10, size: 64$\times$64), room (room-64-64-16, size: 64$\times$64), den312d (den312d, size: 65$\times$81), maze (maze-32-32-4, size: 32$\times$32), and warehouse (warehouse-10-20-10-2-1, size: 161$\times$63).
-Feature All Features SOC Rotations (#) Type-1 edges (#) Type-2 edges (#) Conflict robot pairs (#) Robots (#)
-MAPE 0.0342 0.1778 0.1793 0.2414 0.2826 0.3175 0.4317
-Table 3: Quadratic Regression Results. Except for the first column, each column means that the model is trained with only a single feature
-5.1 Impact of Optimality
-We first use SMART to study the practical effectiveness of the objective used by MAPF algorithms.
-One of the most widely used objectives in MAPF research is the Sum of Costs (SOC), which measures the total number of timesteps for all robots to reach their goals.
-While extensive efforts have been made to develop algorithms that optimize SOC optimally or sub-optimally, it is unclear, for example, whether reducing SOC by 1% translates to a similar degree of improvement in AET in a more realistic simulation.
-In this section, we use SMART to assess the correlation between improvements in SOC and AET.
-Moreover, we want to explore better objectives than SOC in terms of evaluating the quality of the MAPF plan in more realistic settings.
-5.1.1 Experiment Setup
-We generate MAPF plans with varying SOC for a specific MAPF instance by using MAPF-LNS [^3] Code at https://github.com/Jiaoyang-Li/MAPF-LNS Li et al.
-(2021a), an anytime method that iteratively refines MAPF plans to improve SOC within a predefined runtime limit.
-For each map, we select three different numbers of robots and five sets of different start and goal locations for robots.
-For each MAPF instance, we record all plans generated in each iteration of MAPF-LNS within the runtime limit of 60 seconds.
-We then execute these plans in SMART.
-Figure 6: Relationship between SOC and AET. Each dot represents a MAPF plan and its execution result.
-5.1.2 Results and Analysis
-We first study whether SOC can measure the real-world execution time of a MAPF plan.
-As shown in Fig.
-6, SOC and AET show a strong positive correlation across all environments and all numbers of robots, indicating that SOC captures the trend of AET.
-However, the correlation is not perfectly monotonic.
-For the same SOC value, there are noticeable variations in AET, suggesting that SOC alone cannot account for all factors influencing execution time.
-This highlights the limitation of SOC as the sole objective of the MAPF problem.
-To explore other factors that influence execution time, we extract additional features from the ADG, including (1) Type-2 edges (#), the number of type-2 edges, (2) Type-1 edges (#), the number of type-1 edges, (3) Rotations (#), the number of rotation actions, (4) Conflict robot pairs (#), the number of robot pairs connected by type-2 edges, reflecting dependencies in their passing orders, and (5) Robots (#), the number of robots.
-We analyze the correlation between these features and AET by using a quadratic regression model to predict AET based on these features.
-We measure prediction accuracy by using the mean absolute percentage error (MAPE).
-As shown in Table 3, SOC exhibits the highest predictive power among individual features, while other features, such as the number of Type-2 edges, Type-1 edges, and Rotations, also show correlation with AET.
-Meanwhile, the model incorporating all features achieves a much lower MAPE than any single feature, indicating that multiple features influence the execution time.
-In conclusion, incorporating features in addition to SOC is crucial to developing a more accurate objective function for execution time estimation during MAPF planning.
-Figure 7: AET comparison for MAPF with different robot models.
-5.2 Impact of Different Robot Models
-The standard MAPF problem relies on a simplified model that ignores critical execution factors such as kinodynamics and potential delays.
-This simplification creates a gap between planning and real-world execution.
-Thus, recent research has introduced more realistic MAPF models that consider real-world factors.
-An ideal MAPF model should capture the essential aspects of robot kinodynamics while omitting unnecessary details.
-As SMART supports various robot models, we conduct an empirical comparison of different MAPF models to identify the most effective one in real-world scenarios.
-In Varambally et al.
-(2022), a similar problem is studied, but the experiments are limited to a warehouse map and consider no more than 50 robots.
-Additionally, we compare a wider range of robot models.
-5.2.1 Experiment Setup
-We evaluate the impact of incorporating robot kinodynamics and potential delays.
-For the kinodynamics models, we consider: (1) the standard MAPF model, (2) the MAPF model with rotation, and (3) the MAPF model that accounts for the full kinodynamics of robots, including rotation, speed, and acceleration.
-To model delays, we use the $k$-robust delay model Atzmon et al.
-(2020), which ensures that robots can avoid collisions even if delayed by up to $k$ timesteps.
-By combining kinodynamics and delay factors, we evaluate the following MAPF algorithms: (1) MAPF, $k=i$: This method uses the standard MAPF model and a $k$-robust delay model of $k=i$.
-(2) MAPF+rotation, $k=i$: This extends the standard MAPF model by incorporating in-place rotation but does not account for other kinematics.
-It also includes a $k$-robust delay model.
-(3) MAPF+kinodynamics: Derived from Yan and Li (2024a), this method incorporates accurate kinodynamic constraints of robots.
-We use PBS Ma et al.
-(2019) to handle conflicts between robots under these models with a runtime limit of 60 seconds.
-We run these methods to generate MAPF plans for all maps using 25 “random scenarios” with a progressive number of robots until the success rate is smaller than 20%.
-Finally, these plans are executed using SMART.
-Figure 8: Paths generated on different maps.
-5.2.2 Results and Analysis
-As shown in Fig.
-7, the method incorporating the full kinodynamics of robots achieves the shortest AET among all the models.
-Additionally, the planner considering robot rotations demonstrates much better AET than the standard MAPF model.
-These results indicate that incorporating more accurate kinodynamic models into planning can improve execution time.
-However, we also observe trade-offs associated with these models.
-More complex kinodynamic models lead to longer runtimes during planning.
-Meanwhile, the $k$-robust model achieves better or comparable performance with fewer robots with larger values of $k$; however, its performance declines significantly as the number of robots increases, resulting in lower solution quality and reduced scalability.
-Interestingly, in warehouse environments, MAPF+kinodynamics does not provide significant improvements over the MAPF+rotation.
-To better understand this, we visualize the MAPF plans in Fig.
-8, comparing the warehouse map with random and room maps.
-The planned paths show that MAPF+rotation and MAPF+kinodynamics generate paths with more consecutive movements and fewer conflict vertices in warehouse maps than in the other two maps.
-We hypothesize that this minimizes the potential benefits of accounting for kinodynamics in such environments.
-In summary, while more accurate kinodynamic models generally improve execution time in realistic settings, their increased computational complexity can limit scalability.
-Additionally, the $k$-robust model is less efficient and may even harm performance, particularly in more congested scenarios.
-Figure 9: AET using different MAPF models and planners.
-5.3 Optimality and Robot Model
-From Sections 5.1 and 5.2, we show that both improving SOC and incorporating a more accurate kinodynamic model during planning can lead to better AET during execution.
-However, both of them require additional computation time.
-Considering the limited planning time available in real-world applications, this section explores the balance between these factors to provide practical insights.
-5.3.1 Experiment Setup
-We generate MAPF plans with varying SOC for three robot models: (1) standard MAPF, (2) MAPF with rotation, and (3) MAPF with kinodynamics.
-To obtain optimal solutions, we use CBS [^4] Code at https://github.com/Jiaoyang-Li/CBSH2-RTC Li et al.
-(2021b) for the standard MAPF model and CBS with rotation [^5] Code at https://github.com/YueZhang-studyuse/MAPF_T Zhang et al.
-(2023a) for the rotation model.
-For suboptimal plans, we combine PBS with all three models, as described in Section 5.2.
-To analyze the relationship between AET and SOC, we create a naive anytime planner by running PP with random restarts and recording the MAPF plan with the lowest SOC until reaching a runtime limit of 60 seconds.
-We do not use existing anytime solvers, such as MAPF-LNS, because, to our best knowledge, no available implementation considers robot models with kinodynamic constraints.
-5.3.2 Results and Analysis
-As shown in Fig.
-9, planners using more accurate robot models consistently achieve better AET across all maps, even when the methods using less accurate models are optimal.
-This indicates that incorporating more accurate kinodynamic robot models into planning may be more important than obtaining optimal solutions with less accurate models.
-Notably, as discussed in Section 5.1, since SOC is not a perfect metric, the optimal planner does not always achieve the best AET.
-Meanwhile, execution frameworks may also become a bottleneck.
-Since ADG adopts an overly conservative strategy, assuming robots can be delayed indefinitely, it compromises the solution quality of the MAPF plan in certain cases.
-In conclusion, the limiting factors for achieving better AET lie in the more accurate robot models or the execution frameworks with better performance.
-6 Conclusion
-We introduced SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings.
-SMART bridges the gap between methods that use simplified models and their real-world deployments.
+Table II shows the results.
+When using ARGoS3 as the simulator, we observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops.
+The simulation speed in machine (2) is slower than in machine (1) due to an inferior processor.
+In contrast, simulations with Isaac Sim result in significantly higher CPU and memory usage, indicating that Isaac Sim is better suited for scenarios with fewer agents but more powerful hardware.
+Sim Robots CPU Usage Memory Usage Simulation Speed
+(#) (%) (GB)
+[…]
+[…]
+100 16.48 $\pm$ 0.58 0.19 $\pm$ 0.26 33.49 $\pm$ 0.15
+500 22.86 $\pm$ 1.59 0.24 $\pm$ 0.01 13.21 $\pm$ 0.13
+1000 46.02 $\pm$ 0.98 0.65 $\pm$ 0.03 7.37 $\pm$ 0.04
+[…]
+[…]
+100 N/A N/A N/A
+500 N/A N/A N/A
+1000 N/A N/A N/A
+[…]
+50 20.20 $\pm$ 0.91 0.05 $\pm$ 0.03 21.42 $\pm$ 0.35
+100 27.96 $\pm$ 0.39 0.04 $\pm$ 0.00 14.37 $\pm$ 0.10
+500 68.74 $\pm$ 17.24 0.22 $\pm$ 0.01 3.89 $\pm$ 0.02
+1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 1.95 $\pm$ 0.01
+TABLE II: Comparison of CPU usage, memory consumption, and simulation speed for ARGoS3 and Isaac Sim across two machines and various robot counts.
+Since Isaac Sim only supports recent GPUs with ray tracing capabilities, we conducted experiments with Isaac Sim only on Machine (1).
+Figure 8: (a) Experimental setup in Isaac Sim. (b) Experimental setup with real robots.
+V-D Evaluation in Real World
+[…]
+Given a set of start and goal locations, MAPF-LNS2 is used on a central server to generate a MAPF plan.
+SMART then processes this plan and transmits control commands to each robot via a wireless network, with a control frequency of 40 Hz.
+Each JetBot follows its assigned commands accordingly.
+As shown in the accompanying multimedia material, SMART robustly executes the MAPF plan on physical robots, highlighting the practical feasibility and reliability of our planning framework in real-world scenarios.
+Across 10 trials, SMART achieved a 100% success rate, with an average execution time of 48.48 $\pm$ 8.79 seconds and a maximum execution time of 73.83 $\pm$ 9.81 seconds.
+For reference, in Isaac Sim, the average execution time is 45.66 $\pm$ 0.01 seconds, and the maximum execution time is 70.97 $\pm$ 0.00 seconds.
+VI Conclusion
+We introduce SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings.
+SMART bridges the gap between methods that use simplified MAPF models and their real-world deployments.
 SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments.
-Our experiments demonstrated that SMART can handle thousands of robots while maintaining high simulation speeds.
-We highlight several key challenges arising from the application of MAPF in realistic scenarios.
-We further show how SMART can be used to address these challenges.
-Future work includes extending the framework to address lifelong MAPF scenarios, integrating it with other execution frameworks Coskun and O’Kane (2019); Berndt et al.
-(2023); Feng et al.
-(2024), and interfacing it with real robots.
-Acknowledgments
-The research was supported by the National Science Foundation (NSF) under grant number #2328671 and a gift from Amazon.
-The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S.
-government.
+We integrate SMART with two simulators, ARGoS3 and Isaac Sim.
+Our experiments show that SMART can handle thousands of robots while maintaining high simulation speeds when using ARGoS3.
+Future work includes extending the framework to address lifelong MAPF scenarios and integrating it with other execution frameworks [34, 35, 36].

### v3

@@
-Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)
+Advancing MAPF Toward the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)
 Abstract
 We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms.
-MAPF focuses on planning collision-free paths for a group of agents.
-While state-of-the-art MAPF algorithms can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear.
+MAPF focuses on planning collision-free paths for a group of robots.
+While state-of-the-art MAPF planners can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear.
 Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms.
-Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings.
-SMART fills this gap with several advantages: (1) SMART uses physics-engine-based simulators to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots.
-The code is publicly available at https://github.com/smart-mapf/smart.
-Index Terms:
-Multi-Agent Path Finding, Multi-Robot System, Realistic Simulation.
+Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF planners in their specific settings.
+SMART fills this gap with several advantages: (1) SMART uses physics-engine-based simulators to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF planners and robot models, and (3) SMART scales to thousands of robots.
+The code is publicly available at https://github.com/smart-mapf/smart with an online service available at https://smart-mapf.github.io/demo/.
 I Introduction
@@
 In recent years, MAPF has grown into a rapidly expanding research area [19, 32].
-State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds [7, 8, 9].
+State-of-the-art MAPF planners can plan paths for hundreds of robots within seconds [16, 35, 14].
 However, these algorithms make several simplifying assumptions.
@@
 Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors.
-Therefore, while researchers have shown that MAPF algorithms have a wide range of applications, their performance in more realistic settings is unclear.
+Therefore, while researchers have shown that MAPF planners have a wide range of applications, their performance in more realistic settings is unclear.
 Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c, d) Robot models used in the environment. (e) A 3D view of the environment.
-MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings.
-In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF algorithms is impractical.
+MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF planners in realistic settings.
+In this paper, we use the term scalability to refer to runtime capacity in terms of the maximum number of robots that the testbed can support in simulation.
+In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF planners is impractical.
 Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers.
 Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans.
-Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort.
-While some software tools [10, 11] have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps.
+Moreover, since many researchers come from the AI community, learning such robotic simulation tools can require significant effort.
+While some software tools [27, 9] have been developed to evaluate MAPF planners in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF planners and maps.
 In addition, industrial professionals possess realistic simulators for their applications of interest.
-However, they may not have the resources and expertise to implement state-of-the-art MAPF algorithms in their simulators.
-By having a realistic simulator that can easily interface with existing implementations of MAPF algorithms, industrial professionals can more easily understand and select a MAPF algorithm that fits their demands.
-In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios.
-SMART comprises three main components: (1) physics-engine-based simulators, (2) an execution monitoring (EM) server, and (3) robot-specific executors.
+However, they may not have the resources and expertise to implement state-of-the-art MAPF planners in their simulators.
+By having a realistic simulator that can easily interface with existing implementations of MAPF planners, industrial professionals can more easily understand and select a MAPF planner that fits their demands.
+In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF planners in realistic scenarios.
+SMART comprises three main components: (1) a physics-engine-based simulator, (2) an execution monitoring (EM) server, and (3) robot-specific executors.
 The simulator generates realistic simulation environments based on user-defined configurations.
@@
 The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors.
-It leverages the Action Dependency Graph (ADG) [3], a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models.
+It leverages the Action Dependency Graph (ADG) [11], a MAPF execution framework, to ensure robust execution of the paths generated by MAPF planners using various simplified robot models.
 Each robot is assigned an individual executor that receives and executes actions from the server.
@@
 Our contributions in this paper are as follows: 1.
-We present SMART, an open-source software tool for evaluating MAPF algorithms in realistic scenarios.
+We present SMART, an open-source software tool for evaluating MAPF planners in realistic scenarios.
 SMART directly accepts MAPF plans and problem instances from existing benchmarks.
 2.
-We integrate SMART with two simulators, ARGoS3 and Isaac Sim, and robot models and empirically evaluate its scalability and replicability.
-SMART can scale to thousands of robots while maintaining consistent and reproducible results, while prior tools fail to handle a hundred robots.
+We integrate SMART with two simulators, ARGoS3 [25] and Isaac Sim [21], and empirically evaluate its scalability and replicability.
+SMART can scale to thousands of robots while maintaining consistent and reproducible results, whereas prior tools fail to handle a hundred robots.
 3.
@@
 (c) ADG from the MAPF plan.
-Green ovals are actions, with the first element as the reference time from the plan, the second as the action type, and the last two as start and goal locations.
+Green ovals are actions, with the first element as the reference time from the plan, the second as the action type (e.g., rotation or forward movement), and the last two as start and goal locations.
 Horizontal edges are Type-1, and vertical edges are Type-2.
 All robots initially face north.
-Figure 3: System overview of SMART.
 II Background
@@
 Next, we discuss existing execution frameworks for executing MAPF plans.
-Finally, we review benchmarks used to evaluate MAPF algorithms.
+Finally, we review benchmarks used to evaluate MAPF planners.
 II-A MAPF Problem and MAPF Plan
 A MAPF instance [29] consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ constructed from a 2-D grid map and a set of $I$ robots $\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations.
-Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations.
+Given a MAPF instance, a MAPF planner aims to find a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations.
 Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex.
@@
 2 (a) and (b), respectively.
-The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
-Prior research has proposed many MAPF algorithms, most of which are developed based on 2-D grid maps.
+The typical objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
+Prior research has proposed many MAPF planners, most of which were implemented and tested on 2-D grid maps.
 Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees [28, 15], while others focus on scalability, solving MAPF problems involving thousands of robots [22, 23].
-Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency [16, 17].
 Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process [5, 1, 33, 34].
-They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
+They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among robots, while the low level plans for each robot considering real-world factors.
 II-B Executing MAPF Plans
@@
 Thus, we need techniques to execute MAPF plans safely given the uncertainty in execution time.
-One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next.
+One naive way is synchronized execution, where robots need to wait for all robots to finish their current actions before starting the next.
 This ensures strict compliance with the MAPF plan but significantly reduces efficiency.
-Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well.
+Alternatively, one can replan whenever a robot is delayed, but this is computationally expensive and thus does not scale well.
 Therefore, executing MAPF plans using an execution framework is preferred [3].
@@
 The detailed execution procedure is described in Section III-B.
+Although modern MAPF solvers have become significantly faster, execution frameworks such as ADG remain essential due to inevitable real-world uncertainties.
+Some of them, such as communication delays, cannot be eliminated through faster planning alone.
+Moreover, by enforcing only necessary ordering constraints rather than strict timestep synchronization, ADG enables robots to execute multiple consecutive actions without stopping, supporting velocity optimization and smoother execution rather than requiring stop-and-go behavior at every timestep.
+TABLE I: Comparison of SMART with existing multi-robot frameworks and evaluation tools.
+Interactive execution diagnostics refer to inspection of execution states during runtime, and benchmark support refers to compatibility with standard MAPF benchmarks.
+Tool Max Robots MAPF Planner General Execution Interactive Execution Benchmark Primary
+(Tested) Integration Framework Diagnostics Support Purpose
+Open-RMF [24] $\sim$100 Manual Yes No No Fleet deployment
+MRP-Bench [27] <<10 Manual No No Partial Small-scale MAPF evaluation
+REMROC [9] <<50 Manual No No No Human-aware evaluation
+SMART (ours) 2000 Yes Yes Yes Yes MAPF research
 II-C Evaluating MAPF Plans
-To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it.
-While many MAPF benchmarks with 2-D-grid maps are proposed [1, 24, 25, 26], most prior MAPF works are evaluated on them using simplified robot models that assume no delays and discretized time and space.
+To evaluate MAPF planners, researchers typically generate a plan for a given MAPF instance and then execute it.
+While many MAPF benchmarks with 2-D grid maps are proposed [29, 8, 30, 26], most prior MAPF works are evaluated on them using simplified robot models that assume no delays and discretized time and space.
 Some benchmarks have attempted to include realistic elements such as acceleration and rotation [20, 4].
-However, their model is still far from realistic, and typically requires the same robot models for both planning and execution phases.
+However, their models are still far from realistic, and they typically require the same robot models for planning and execution phases.
 Given the difficulty of modeling all real-world factors analytically, physics-engine-based simulators offer a promising approach for more accurate evaluation.
+Frameworks like Open-RMF [24] provide middleware for deploying and coordinating robot fleets in production environments, handling task allocation, traffic management, and building infrastructure integration.
+However, such operational deployment frameworks are not specialized for MAPF evaluation.
+While centralized MAPF planners can be integrated, using them to benchmark and stress-test MAPF planners at large scale typically requires additional engineering effort beyond their core scope.
+To address the need for MAPF-specific evaluation, several simulation-based tools have been developed.
 MRP-Bench [27] leverages a high-fidelity, physics-based simulator Gazebo [13] to simulate multi-robot systems.
@@
 REMROC [9] also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments.
-However, these existing tools have two main limitations.
-First, they often require significant engineering effort to integrate user-defined maps and MAPF planners, limiting their flexibility for different scenarios.
-In particular, they lack a general execution framework, making it difficult to directly evaluate MAPF planners using different robot models.
-Second, they have limited scalability: the largest number of robots tested in previous work is less than 100 agents.
+However, as shown in Table I, these existing tools have two main limitations.
+First, they often require significant engineering effort to integrate user-defined maps and new MAPF planners, limiting their flexibility for different scenarios.
+In particular, they lack a general execution framework, making it difficult to directly evaluate MAPF planners across different robot models assumed during planning.
+Second, they have limited scalability: the largest number of robots tested in previous work is less than 100 robots.
 Most of these works rely on Gazebo, whose high computational overhead, resulting from modeling many factors not relevant to MAPF, significantly constrains scalability.
 MRP-Bench, for example, does not scale to more than 10 robots.
+Figure 3: System overview of SMART.
 III SMART Architecture
@@
 III-A Simulator
-Given a user-provided configuration, the simulator creates realistic simulation environments.
-The configuration specifies obstacle locations and shapes, robot start locations, and detailed robot configurations (e.g., the robot model, as well as its velocity and acceleration limits).
-To support the widely used MovingAI benchmark [1], SMART can converts its 2-D grid maps and corresponding MAPF instances into the configuration required by the simulator.
-During execution, the simulator uses a physics engine to update the states of the robots and environment, including their poses, dynamics, sensors, and actuators, based on the received control commands from the executors.
-The executors can query the simulator for real-time robot information, such as their current pose.
-To accommodate diverse user requirements, SMART supports two simulation platforms: ARGoS3 [30] and Isaac Sim [31].
+Given a user-provided configuration, the simulator creates a realistic simulation environment.
+The configuration specifies obstacle locations and shapes, robot start locations, and robot configurations (e.g., velocity and acceleration limits).
+To support the widely used MovingAI benchmark [29], SMART can convert its 2-D grid maps and corresponding MAPF instances into the configuration required by the simulator.
+During execution, the simulator uses a physics engine to update the states of the robots, including their poses, dynamics, and actuators, based on the received control commands from the executors.
+The simulator provides real-time robot state information (e.g., current pose) to other system components.
+To accommodate diverse user requirements and balance scalability with physical realism, SMART supports two simulation platforms: ARGoS3 [25] for efficient large-scale experiments and Isaac Sim [21] for higher-fidelity simulation.
 III-A1 ARGoS3
@@
 1 (d)), a differential-drive robot commonly used in research and education.
-Figure 4: (a) Cycle in the ADG causing all robots to be stalled. (b) Adding safe stops between vertices to make the ADG acyclic.
+Figure 4: (a) Cycle in the ADG causing all robots to be stalled.
+(b) Adding safe stops between vertices to make the ADG acyclic.
+Robot diameter must be at most half the cell size to prevent collisions at safe stops between vertices.
 III-B Execution Monitoring (EM) Server
@@
 Server Initialization
-During the initialization, we use the MAPF plan parser to construct an ADG.
-We define a standardized format to store the MAPF plan as a sequence of timed locations for each robot $r_{i}$, i.e., $((x_{i}^{0},y_{i}^{0},t_{i}^{0}),...,(x_{i}^{T},y_{i}^{T},t_{i}^{T}))$.
-The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans.
-This allows SMART to work with plans using varying simplified robot models.
+During initialization, we use the MAPF plan parser to convert it into a sequence of actions and then construct an ADG based on these actions.
+We define a lightweight, planner-agnostic standardized format to store the MAPF plan as a sequence of timed locations for each robot $r_{i}$, i.e., $((x_{i}^{0},y_{i}^{0},t_{i}^{0}),...,(x_{i}^{T},y_{i}^{T},t_{i}^{T}))$.
+This simple representation can be easily derived from any MAPF-generated path, making the parser compatible with any valid MAPF plan regardless of the underlying robot model or planner used to generate them.
+This design choice allows SMART to work seamlessly with planners using varying simplified robot models without requiring modifications to the planners themselves.
 The parser first converts the path of each robot into a sequence of actions.
-Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path and adds a rotation action when needed.
-The ADG is constructed based on the sequence of actions.
-ADG requires the MAPF plan to be circle-conflict free, ensuring that robots do not collide even if delayed by one timestep [3].
+Specifically, for each transition between consecutive vertices, the parser infers the required robot orientation.
+If the orientation differs from the current orientation, an in-place rotation action is added.
+Subsequently, a forward movement action is added to move the robot to the next vertex.
+The ADG is then constructed based on this sequence of actions.
+ADG requires the MAPF plan to be cycle-conflict-free, i.e., it prohibits cyclic location exchanges among a set of robots within the same timestep.
+This is because ADG only allows movements into empty space to ensure delay safety, which cannot be satisfied under a cycle conflict [11].
 However, ensuring this is not standard in most MAPF models.
-As a result, if we directly build an ADG based on a plan with circle conflict, we may create deadlocks, such as the one illustrated in Fig.
+As a result, if we directly build an ADG based on a plan with cycle conflicts, we may create deadlocks, such as the one illustrated in Fig.
 4 (a).
@@
 4 (b), making the ADG compatible with these MAPF models.
-Execution Managing
+Execution Management
 During execution, an execution monitor framework manages the progress of execution and communicates with robot executors.
@@
 A vertex is marked as finished when the executor confirms the completion of its associated action with the server.
-Communication with robot executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
+Communication with executors is managed by a Remote Procedure Call (RPC) server built with rpclib.
+[^1] Available at http://rpclib.net.
 III-C Executor
@@
 3, each executor consists of an RPC client, an action queue, and a controller.
-The RPC client establishes communication with the server, receiving enqueued actions and sending confirmation for finished actions.
-The action queue keeps track of the enqueued actions, which are executed in a first-in-first-out manner.
-A Proportional Integral Derivative (PID) controller is used to execute these actions by sending control commands to the simulator.
-Once an action is finished, a confirmation message is sent to the server.
+The RPC client communicates with the server by receiving enqueued actions and sending confirmations for finished actions.
+The action queue stores enqueued actions and executes them in first-in-first-out order.
+A Proportional Integral Derivative (PID) controller executes these actions by sending control commands to the simulator, and a confirmation is sent to the server upon completion.
 To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue.
 For example, if a robot receives three consecutive forward movement actions, the executor merges them into a single longer forward action, reducing unnecessary stop-and-go behavior.
+This distributed executor design enables parallel operation, as each robot handles the above-mentioned operations independently, improving scalability.
 III-D Performance Metrics
@@
 SMART also provides visualization for the MAPF plan executions, allowing researchers to better understand the behaviors of the robots.
-IV Using SMART
-Users can evaluate their solutions using SMART in two ways: by uploading their MAPF problem and solution through a web interface or by deploying SMART locally.
+IV System Usage
+Users can evaluate their solutions using SMART in two ways: by uploading their MAPF problem instance and plan through a web interface or by deploying SMART locally.
 Figure 5: The SMART web interface visualizing a multi-robot simulation on a grid map.
 In the viewport, users interact with individual robots to inspect paths and inter-robot constraints.
-On the side panel, users can configure the environment, control playback, and see execution details.
+On the side panel, users can configure the environment, robot kinematic limits, control playback, and see execution details.
 SMART as a Service
-SMART, with ARGoS3 and the 3D visualizer, is directly available as software-as-a-service [^2] The link will be made available here upon acceptance..
+SMART, with ARGoS3 and the 3D visualizer, is directly available as software-as-a-service.
 Specifically, we provide an intuitive web app for running and inspecting MAPF simulations.
-Complete with a beautiful 3D-visualizer built with React and Three.js, it allows researchers to instantly understand how algorithms perform with physical robots (Fig.
+It features a 3D visualizer built with React and Three.js, allowing researchers to better understand how algorithms perform on physical robots (Fig.
 5).
 Users configure the environment by specifying a map and scenario file in the MovingAI Benchmark format.
-Within the online interface, after users upload the files and select the desired kinodynamic settings for the robots, and initiate the simulation with a single click.
+Within the online interface, after users upload the files and select the desired kinodynamic settings for the robots, users can initiate the simulation with a single click.
 After starting the simulation, they can control the simulation time step and adjust the playback speed.
-Users may select robots to view complete paths, execution status, and Type-2 edges between robots.
+Users may select robots to view full paths, execution status, and Type-2 edges between robots.
 Installing SMART Locally
@@
 In ARGoS3, each grid cell represents 1 meter.
-Each foot-bot robot is subject to a speed limit of [0, 5] m/s, an acceleration limit of [-0.4, 0.4] m/s2, and an angular velocity limit of 30/s.
+Each foot-bot robot is subject to speed limits of [0, 5] m/s, acceleration limits of [-0.4, 0.4] m/s2, and absolute angular velocity limits of 30/s.
 In Isaac Sim, each grid cell represents 0.5 meters.
-Each JetBot is constrained to a speed limit of [0, 2.5] m/s, an acceleration limit of [-0.2, 0.2] m/s2, and the same angular velocity limit of 30/s.
-We evaluated our system using MAPF plans generated by three different methods: CBS [12], PBS [7], and MAPF-LNS2 [9].
+Each JetBot is limited to a speed limit of [0, 2.5] m/s, an acceleration limit of [-0.2, 0.2] m/s2, and the angular velocity limit of 30/s.
+We evaluate SMART using MAPF plans generated by MAPF-LNS2 [14], which provides strong scalability and is representative of state-of-the-art MAPF solvers.
+Unless otherwise specified, all results in this section are based on MAPF-LNS2.
+To verify generality, we additionally conducted experiments using plans generated by other MAPF planners, including CBS [28] and PBS [16].
 These plans were created using four different robot models: the standard MAPF model [29], the $k$-robust delay model [1], MAPF with rotation [36], and MAPF with kinodynamics [34].
-Across all combinations, our system achieved a 100% execution success rate.
-In the remainder of this section, we present results based on MAPF plans generated by MAPF-LNS2, as it demonstrated the highest scalability among the algorithms evaluated.
+Across all tested settings, SMART achieved a 100% execution success rate, demonstrating its compatibility with different MAPF planners.
 V-A Scalability and Runtime Performance
@@
 For example, if the simulation speed is 10, the simulation world runs 10 times faster than in real time.
-MAPF plans are generated using MAPF-LNS2333Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 [9] with a progressively increasing number of robots and a time limit of 60 seconds.
+MAPF plans are generated using MAPF-LNS2222Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 [14] with a progressively increasing number of robots and a time limit of 60 seconds.
 For each map, we randomly select 5 scenarios from the MovingAI Benchmark to generate plans, which are then executed in SMART without visualization.
@@
 In contrast, for Isaac Sim, the upper bound on the number of robots is determined by the point at which the simulation speed drops below real-time (i.e., simulation speed << 1).
+Figure 7: Simulation speed in all environments.
+TABLE II: Replicability using different numbers of robots.
 Simulator Robots average execution maximum execution
@@
 50 183.71 $\pm$ 0.07 435.46 $\pm$ 1.36
-TABLE I: Replicability using different numbers of robots.
-Figure 7: Simulation speed on all environments.
 As shown in Fig.
 7, SMART supports simulations with up to 2,000 robots.
-The more robots and the larger the map, the smaller the simulation speed.
+As the number of robots increases and the map becomes larger, the simulation speed decreases.
 Nevertheless, for ARGoS3, SMART maintains a simulation speed of around 10.0 with 1,000 robots in warehouse.
@@
 In contrast, Isaac Sim provides the capability to model more complex real-world factors while also producing photorealistic visualizations, but it is significantly less scalable, particularly in large environments.
-It is therefore more suitable for experiments with a smaller number of agents.
+It is therefore more suitable for experiments with a smaller number of robots.
 V-B Replicability
@@
 For each run, we record both the average and the maximum execution time.
-As shown in Table I, SMART demonstrates strong replicability across runs, with minor execution time variations.
+As shown in Table II, SMART demonstrates strong replicability across runs, with minor execution time variations.
 While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
@@
 We use machine (2) to demonstrate that SMART can run on resource-constrained machines.
-To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark [1].
+To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MovingAI Benchmark [29].
 All simulations are parallelized on all CPU cores on both machines.
 We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots.
-Table II shows the results.
+Table III shows the results.
 When using ARGoS3 as the simulator, we observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops.
 The simulation speed in machine (2) is slower than in machine (1) due to an inferior processor.
-In contrast, simulations with Isaac Sim result in significantly higher CPU and memory usage, indicating that Isaac Sim is better suited for scenarios with fewer agents but more powerful hardware.
+In contrast, simulations with Isaac Sim result in significantly higher CPU and memory usage, indicating that Isaac Sim is better suited for scenarios with fewer robots but more powerful hardware.
+TABLE III: Comparison of CPU usage, memory consumption, and simulation speed for ARGoS3 and Isaac Sim across two machines and various robot counts.
+Since Isaac Sim only supports recent GPUs with ray tracing capabilities, we conduct experiments with Isaac Sim only on Machine (1).
 Sim Robots CPU Usage Memory Usage Simulation Speed
@@
 1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 1.95 $\pm$ 0.01
-TABLE II: Comparison of CPU usage, memory consumption, and simulation speed for ARGoS3 and Isaac Sim across two machines and various robot counts.
-Since Isaac Sim only supports recent GPUs with ray tracing capabilities, we conducted experiments with Isaac Sim only on Machine (1).
 Figure 8: (a) Experimental setup in Isaac Sim. (b) Experimental setup with real robots.
@@
 VI Conclusion
-We introduce SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings.
+We introduce SMART, a scalable testbed for evaluating MAPF planners under realistic settings.
 SMART bridges the gap between methods that use simplified MAPF models and their real-world deployments.
@@
 Our experiments show that SMART can handle thousands of robots while maintaining high simulation speeds when using ARGoS3.
-Future work includes extending the framework to address lifelong MAPF scenarios and integrating it with other execution frameworks [34, 35, 36].
+SMART’s modular, planner-agnostic design supports diverse application-driven use cases across academic and industrial settings.
+Academic researchers can easily benchmark different algorithms without custom integration, while industrial users can evaluate algorithms in their specific warehouse or manufacturing environments without requiring specialized MAPF expertise.
+Future work includes: (1) extending to lifelong MAPF [17] where tasks arrive continuously and require dynamic replanning, (2) integrating sophisticated dynamic obstacle models for human-robot interaction scenarios, (3) supporting alternative execution frameworks [6, 2, 7].
+Acknowledgments
+The research was supported by the National Science Foundation (NSF) under grant numbers #2328671 and #2441629, as well as a gift from Amazon.
+The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S.
+government.